# How hydrodynamic interactions alter polymer stretching in turbulence

### v1

@@ II.2 Brownian dynamics simulations
From Eq. (1), we see that the centre of mass of the chain moves like a tracer in the turbulent flow, since its Brownian motion is much weaker than turbulent advection. Further, as we are focusing on the the ultra-dilute limit in which the flow remains unaffected by elastic feedback forces, the tracer-trajectory of a chain’s centre of mass is independent of its internal structure. Therefore, Eq. (2) for the chains extension may be solved independently, given the Lagrangian data of the velocity gradient $\bm{\kappa}(t)$ along a tracer trajectory. Here, we use pre-stored Lagrangian data from a direct numerical simulation (DNS) of homogeneous isotropic incompressible turbulence, as well as a random Gaussian model for $\bm{\kappa}(t)$ (see Sec. II.3).
@@
Equation (2) is integrated using the Euler-Maruyama method. The rejection algorithm of [49] is used to prevent $Q$ from exceeding $Q_{\mathrm{m}}$ and causing a numerical divergence of the FENE force. Any time-step update that yields extensions greater than $Q_{\mathrm{m}}\big{(}1-\sqrt{\delta t/10}\big{)}^{1/2}$ is rejected. By using sufficiently small time steps $\delta t$, we have limited such rejections to a negligible fraction of the total steps in a simulation. Note that, unlike the case of extensional flow, polymers in turbulence do not persistently align with the straining direction of the velocity gradient due to constant vorticity-induced rotation. Hence, sophisticated implicit time-integration methods are not needed here. As a check, though, we have redone a few simulations, with and without HI, using the semi-implicit, predictor-corrector adaptive time-stepping algorithm of Prabhakar and Prakash [9], and find that our results for the PDF of extension remain unchanged.
@@
Each chain starts out with a random configuration and is allowed to equilibrate in a still fluid ($\bm{\kappa}=0$). The flow is then ’turned on’, with $\bm{\kappa}$ drawing values from either the DNS dataset or the random model. Since the distribution of chains starts out from an equilbrium distribution, some time is required for the chains to stretch out in the fluctuating flow and for the stationary distribution of extension to be established. In this work, we ignore the transients and focus on the stationary statistics.
II.3 Velocity gradients in turbulent and random flows
@@
To study polymer stretching in homogeneous isotropic turbulence, we use a database of velocity gradients $\bm{\kappa}(t)$, stored along the trajectories of $2\times 10^{5}$ tracers. The corresponding DNS, with Taylor-microscale Reynolds number $\mathit{Re}_{\lambda}\approx 111$, was performed in a tri-periodic cube, wherein the incompressible Navier–Stokes equations were solved using a standard fully-dealiased pseudo-spectral method with $512^{3}$ grid points [55]. Time integration was performed using a second-order Adams–Bashforth scheme. The tracers were evolved in the flow using a second-order Runge–Kutta method [55]. We have calculated the Lyapunov exponent $\lambda$ along these tracer trajectories from $\bm{\kappa}(t)$ using the continuous $QR$ method [56]. Further details on the DNS data set are provided in Appendix A.
@@
II.4 Dependence of chain parameters on the number of beads
To understand the effect of HI on chains with different levels of coarse-graining, we need a way to compare the results of chains with different numbers of beads. In the absence of HI, the mapping of Jin and Collins [58] has been found to work very well in homogeneous isotropic turbulent flows. Specifically, the PDF of $R/R_{\mathrm{eq}}$ of two chains with different values of $N_{\mathrm{b}}$ are found to be in agreement, provided the parameters are determined in the following manner. First, we fix the equilibrium length of the links $Q_{\mathrm{eq}}$. This choice also fixes the equilibrium length of the dumbbell version of the model, $R_{\mathrm{eq}}^{D}=Q_{\mathrm{eq}}$. Next, we choose values for the maximum extension and the elastic time scale for the dumbbell model: $R_{\mathrm{m}}^{D}$ and $\tau^{D}$. We then calculate the remaining link parameters, $Q_{\mathrm{m}}$ and $\tau_{s}$, for a chain with $N_{\mathrm{b}}$ beads according to
$\frac{Q_{\mathrm{m}}}{Q_{\mathrm{eq}}}=\frac{1}{\sqrt{N_{\mathrm{b}}-1}}\,\frac{R_{\mathrm{m}}^{D}}{R_{\mathrm{eq}}^{D}},\quad\tau_{s}=\frac{\quad 6\tau^{D}}{(N_{\mathrm{b}}+1)N_{\mathrm{b}}}.$ (8)
This mapping is consistent with the view that the chain is a coarse-grained representation of a polymer with a fixed number of Kuhn steps $N_{\mathrm{k}}$. Recalling that ${N_{\mathrm{k}}}^{1/2}$ is proportional to the ratio of the contour length of the chain $R_{\mathrm{m}}$ to its equilibrium r.m.s. extension $R_{\mathrm{eq}}$, we see that maintaining $N_{\mathrm{k}}$ constant as we coarse-grain requires $R_{\mathrm{m}}/R_{\mathrm{eq}}$ to be constant. Now, for a fixed $Q_{\mathrm{eq}}$, we have $R_{\mathrm{eq}}=\sqrt{N_{\mathrm{b}}-1}Q_{\mathrm{eq}}$ (based on the random-walk theory for a Rouse chain [59, 48]). So, the mapping (8) yields $R_{\mathrm{m}}/R_{\mathrm{eq}}=(N_{\mathrm{b}}-1)Q_{\mathrm{m}}/\sqrt{N_{\mathrm{b}}-1}Q_{\mathrm{eq}}=R_{\mathrm{m}}^{D}/Q_{\mathrm{eq}}$, which is independent of $N_{\mathrm{b}}$ as we have fixed the values of both $R_{\mathrm{m}}^{D}$ and $Q_{\mathrm{eq}}$. Thus, the Jin-Collins mapping sets the link parameters so that the chain’s contour length $R_{\mathrm{m}}$ increases as $\sqrt{N_{\mathrm{b}}-1}$, thereby ensuring that $R_{\mathrm{m}}/R_{\mathrm{eq}}$ is independent of $N_{\mathrm{b}}$. Since $R_{\mathrm{m}}/R_{\mathrm{eq}}=R_{\mathrm{m}}^{D}/Q_{\mathrm{eq}}$, we see that the initial choices for $R_{\mathrm{m}}^{D}$ and $Q_{\mathrm{eq}}$ sets the extensibility of the polymer for all levels of coarse graining, including that of the dumbbell (for which $R_{\mathrm{eq}}^{D}=Q_{\mathrm{eq}})$. In our simulations, we set $Q_{\mathrm{eq}}=1$ and $R_{\mathrm{m}}^{D}=109.54$, yielding an extensibility ratio of ${N_{\mathrm{k}}}=R_{\mathrm{m}}^{2}/R_{\mathrm{eq}}^{2}=12000$. While being relevant to long DNA molecules [19], which have been used in drag-reduction experiments [60, 61], this value of the extensibility ratio ensures a sufficient range of extension for a power-law distribution to manifest.
The time-scale $\tau^{D}$ in (8) may be thought of as an approximation to the longest relaxation time of the entire chain. Its relation to the time-scale of the individual springs $\tau_{s}$ is consistent with the behavior of the largest relaxation time of a Rouse chain (no HI), which for large $N_{\mathrm{b}}$ scales as $N_{\mathrm{b}}^{2}\tau_{s}$. The precise form of the relation between $\tau^{D}$ and $\tau^{s}$ was obtained by Jin and Collins [58] on substituting the relation between $R_{\mathrm{eq}}$ and $Q_{\mathrm{eq}}$ into an expression for the elongational viscosity of highly-stretched polymers, given by Wiest and Tanner [62], and requiring the resulting viscosity values to be independent of $N_{\mathrm{b}}$.
In the absence of HI, the Jin-Collins mapping has been found to work very well for bead-spring chains in homogeneous isotropic turbulent flows: on using the mapping, the PDF of $R/R_{\mathrm{eq}}$ of a dumbbell matches that of a chain with 10 or 20 beads. This is true, not only for the stationary PDF [45], but also for the evolving PDF that describes the initial stretching of polymers when they are first introduced into a turbulent flow [39]. Here, we use the Jin-Collins mapping as a basis of comparison that is particularly well-suited to revealing the effects of HI, given that it fully accounts for changes due to varying $N_{\mathrm{b}}$ in the absence of HI.
We perform simulations of chains with $N_{\mathrm{b}}$ equal to 2 (dumbbells), 4, 10, and 20. The computational cost of performing Brownian dynamics simulations, over a large ensemble of trajectories in the turbulent flow, limits us to twenty-bead chains. We consider a range of values of the Weissenberg number, defined as $\mathrm{Wi}=\lambda\tau^{D}$. The Lyapunov exponent $\lambda$, which is computed along tracer trajectories in the turbulent carrier flow (Sec. II.3), characterizes the long-time asymptotic exponential-stretching behaviour of line elements. Therefore $\lambda^{-1}$ is the most suitable flow time-scale for defining $\mathrm{Wi}$. Indeed, with this definition, the distribution of polymer extensions undergoes a coil-stretch transition near $\mathrm{Wi}\approx 1/2$, as was first anticipated theoretically by Balkovsky et al. [34, 35] and Chertkov [36], and later demonstrated for free-draining dumbbells and chains in homogeneous isotropic turbulence [41, 39].
@@ II.4 Dependence of chain parameters on the number of beads
For a given polymer and solvent, the value of $\mathrm{Wi}$ can be increased by intensifying the turbulence, so that $\lambda$ increases. Consider a stationary homogeneous isotropic turbulent flow that is driven by a large-scale forcing, which injects energy at a rate $\epsilon$. The turbulent properties of the flow are entirely characterized by the Reynolds number $Re=UL\rho/\mu$, or alternatively by the Taylor Reynolds number $Re_{\lambda}\sim Re^{1/2}$. Here, $U$ is the r.m.s. velocity, $L$ is the size of the largest eddies, and $\rho$ is the density of the fluid. Now, $\lambda$ is proportional to the inverse of the Kolmogorov time-scale $\tau_{\eta}$ [63], which is the turnover time of the smallest eddies in the flow. Using the definition $\tau_{\eta}=(\mu/\rho\epsilon)^{1/2}$, and the inertial-range scaling relation $\epsilon\sim U^{3}/L$, one has $\tau_{\eta}=(\nu L)^{1/2}U^{-3/2}=(L/U)Re_{\lambda}^{-1}$ [64]. Usually, $L$ is set by the geometry of the system or the forcing mechanism, so that as one increases the forcing strength, $U$ will increase while $L$ will not. As a result, $\tau_{\eta}$ will decrease and $\lambda$ and $\mathrm{Wi}$ will increase; in addition, $Re_{\lambda}$ will increase. At first glance, the latter variation of $Re_{\lambda}$ seems incompatible with our simulations, which are based on a database of turbulent velocity gradients that are computed at a single fixed value of $Re_{\lambda}$ (see Sec. II.3). One can, in principle, increase $\mathrm{Wi}$ while keeping $Re_{\lambda}$ constant, by adjusting the forcing so that $U$ increases while $L$ decreases (leaving $UL$ unchanged). However, we expect our results to hold to a good approximation even if $Re_{\lambda}$ increases along with $\mathrm{Wi}$. As $Re_{\lambda}$ increases, the distribution of strain-rates will become more heavy-tailed, i.e., the polymer will experience more extreme straining events [65, 66]. However, non-Gaussian extreme strain-rates have only a weak effect on polymer stretching; this result was shown for Rouse chains by [39] and is verified for chains with HI in Sec. V. So, we expect our results on the effect of increasing $\mathrm{Wi}$ to be relevant even to situations in which the increase of $\mathrm{Wi}$ is accompanied by an increase in $Re_{\lambda}$.
The value of $h^{\star}$ is, in principle, tied to the value of $\tau_{s}$ through the bead radius $a$ [see (6) and recall that $\tau_{s}=\zeta/4H$ where $\zeta=6\pi\mu a$]. In practice, however, this relationship is often ignored and $h^{\star}$ is treated as a free parameter that controls the strength of HI [67, 11]; typically, its value is tuned to reproduce experimental observations, e.g., of the the longest relaxation time [67]. Here, we vary $h^{\star}$ freely in order to understand the influence of HI on the stretching dynamics of the chain; specifically, we use $h^{\star}=0$ (free-draining), $h^{\star}=0.2$ (an intermediate and typical value [8, 9]), and $h^{\star}=0.49$ (the largest physically meaningful value).
@@ III Dumbbells and chains respond differently to HI
Figure 1: Mean extension $\langle R\rangle/R_{eq}$ for various values of the Weissenberg number $\mathrm{Wi}$, considering $h^{\star}$ = 0, 0.2, 0.49 [see the legend in panel (a)]. Results are presented for (a) dumbbells, (b) twenty-bead chains. (c) four-bead chains, and (d) ten-bead chains. The inset in panel (a) is a zoom of the main panel. The three curves in panel (c) almost entirely overlap.
On comparing the results without HI ($h^{\star}=0$) to those with HI ($h^{\star}=0.2$ and $0.49$), we see that HI marginally increases the stretching of dumbbells [Fig. 1(a)], while it significantly decreases the stretching of twenty-bead chains [Fig. 1(b)]. The effect of HI changes gradually from weak promotion to strong suppression of extension, as $N_{\mathrm{b}}$ is increased from 2 to 20, as evidenced by the results for $N_{\mathrm{b}}$ = 4 and 10 in Fig. 1(c) and Fig. 1(d), respectively. The slight enhancement of extension in case of dumbbells is in agreement with recent analytical results, derived using the Batchelor-Kraichnan model of turbulent transport [24].
@@
So why do HI affect dumbbells and multi-bead chains differently? At small $\mathrm{Wi}$, chains retract into coiled configurations in which the outer beads hydrodynamically shield the inner beads and thereby weaken the effect of fluid drag on the chain. As a consequence, small $\mathrm{Wi}$ chains stretch less in the presence of HI. Now, a dumbbell cannot form a coil and is by definition always straight. In this constrained arrangement, the effect of the disturbance flow created by the two beads is to increase the apparent relaxation time scale (as is analytically demonstrable [24] via the consistent-preaveraging approximation [68]). To check whether the influence of HI on multibead chains is associated with the formation of physical coils, we measure the hinge angles $\theta_{i}$ ($i=1,2,\dots,N_{\mathrm{b}}-2$) between successive links, and then compute the average angle $\overline{\theta}=\left(\sum_{1}^{N_{\mathrm{b}}-2}\theta_{i}\right)/(N_{\mathrm{b}}-2)$ for each chain. The PDF of $\overline{\theta}$ is presented for twenty-bead chains, for $\mathrm{Wi}=0.3$ and $0.5$, in Figs. 2(a-b). Increasing $h^{\star}$ increases the proportion of configurations with average inter-link angles close to $90^{\circ}$, which is consistent with HI acting to maintain chains in coiled configurations.

### v2

@@ II.3 Brownian dynamics simulations
From Eq. (1), we see that the centre of mass of the chain moves like a tracer in the turbulent flow, since its Brownian motion is much weaker than turbulent advection. Further, as we are focusing on the the ultra-dilute limit in which the flow remains unaffected by elastic feedback forces, the tracer-trajectory of a chain’s centre of mass is independent of its internal structure. Therefore, Eq. (2) for the chains extension may be solved independently, given the Lagrangian data of the velocity gradient $\bm{\kappa}(t)$ along a tracer trajectory. Here, we use pre-stored Lagrangian data from a direct numerical simulation (DNS) of homogeneous isotropic incompressible turbulence, as well as a random Gaussian model for $\bm{\kappa}(t)$ (see Sec. II.4).
@@
Equation (2) is integrated using the Euler-Maruyama method. The rejection algorithm of Ref. [54] is used to prevent $Q$ from exceeding $Q_{\mathrm{m}}$ and causing a numerical divergence of the FENE force. Any time-step update that yields extensions greater than $Q_{\mathrm{m}}\big(1-\sqrt{\delta t/\tau_{\mathrm{s}}}\big)^{1/2}$ is rejected [54]. By using sufficiently small time steps $\delta t$, we have limited such rejections to a negligible fraction of the total steps in a simulation. Note that, unlike the case of extensional flow, polymers in turbulence do not persistently align with the straining direction of the velocity gradient due to constant vorticity-induced rotation. Hence, sophisticated implicit time-integration methods are not needed here. As a check, though, we have redone a few simulations, with and without HI, using the semi-implicit, predictor-corrector adaptive time-stepping algorithm of Prabhakar and Prakash [9], and find that our results for the PDF of extension remain unchanged.
We perform simulations of chains with $N_{\mathrm{b}}$ equal to 2 (dumbbells), 4, 10, and 20. The computational cost of performing Brownian dynamics simulations, over a large ensemble of trajectories in the turbulent flow, limits us to twenty-bead chains.
We consider a range of values of the Weissenberg number, defined as $\mathrm{Wi}=\lambda\tau^{D}$. The Lyapunov exponent $\lambda$, which is computed along tracer trajectories in the turbulent carrier flow (Sec. II.4), characterizes the long-time asymptotic exponential-stretching behaviour of line elements. Therefore $\lambda^{-1}$ is the most suitable flow time-scale for defining $\mathrm{Wi}$. Indeed, with this definition, the distribution of polymer extensions undergoes a coil-stretch transition near $\mathrm{Wi}\approx 1/2$, as was first anticipated theoretically by Balkovsky et al. [37, 38] and Chertkov [39], and later demonstrated for free-draining dumbbells and chains in homogeneous isotropic turbulence [44, 42].
@@
For every combination of parameter values we perform Brownian dynamics simulations of $2\times 10^{5}$ independent trajectories, each of temporal length $75\lambda^{-1}$ (longer computations up to $300\lambda^{-1}$ are performed for the persistence-time analysis of Sec. IV.3). Each chain starts out with a random configuration and is allowed to equilibrate in a still fluid ($\bm{\kappa}=0$). The flow is then ’turned on’, with $\bm{\kappa}$ drawing values from either the DNS dataset or the random model. Since the distribution of chains starts out from an equilbrium distribution, some time is required for the chains to stretch out in the fluctuating flow and for the stationary distribution of extension to be established. In this work, we ignore the transients and focus on the stationary statistics. The time step of the simulations is varied according to $\mathrm{Wi}$ so that $\delta t\approx 0.05\tau_{\mathrm{s}}$ (e.g., for $\mathrm{Wi}=1$, $\delta t=7.5\times 10^{-4}\lambda^{-1}=5.5\times 10^{-3}\tau_{\eta}=0.05\tau_{\mathrm{s}}$).
II.4 Velocity gradients in turbulent and random flows
@@
To study polymer stretching in homogeneous isotropic turbulence, we use a database of velocity gradients $\bm{\kappa}(t)$, stored along the trajectories of $2\times 10^{5}$ tracers. The corresponding DNS, with Taylor-microscale Reynolds number $\mathit{Re}_{\lambda}\approx 111$, was performed in a tri-periodic cube, wherein the incompressible Navier–Stokes equations were solved using a standard fully-dealiased pseudo-spectral method with $512^{3}$ grid points [66]. Time integration was performed using a second-order Adams–Bashforth scheme. The flow was driven to a statistically stationary state via time-dependent large-scale forcing that injected energy at a fixed rate into the first and second wavenumber shells. The corresponding Fourier modes of the instantaneous velocity field were summed and then multiplied by a time-dependent factor to yield the forcing function; the instantaneous value of the factor is determined by the ratio of the desired energy injection rate and the instantaneous kinetic energy content of the first two wavenumber shells. The tracers were evolved in the stationary flow using a second-order Runge–Kutta method [66]. We have calculated the Lyapunov exponent $\lambda$ along these tracer trajectories from $\bm{\kappa}(t)$, using the continuous $QR$ method [67, 68, 69]; in terms of the Kolmogorov time-scale, we obtain $\lambda=0.136/\tau_{\eta}$ in keeping with previous simulations of isotropic turbulence [70]. Further details on the DNS data set are provided in Appendix A.
@@
For a given polymer and solvent, the value of $\mathrm{Wi}$ can be increased by intensifying the turbulence, so that $\lambda$ increases. Consider a stationary homogeneous isotropic turbulent flow that is driven by a large-scale forcing, which injects energy at a rate $\epsilon$. The turbulent properties of the flow are entirely characterized by the Reynolds number $Re=UL\rho/\mu$, or alternatively by the Taylor Reynolds number $Re_{\lambda}\sim Re^{1/2}$. Here, $U$ is the r.m.s. velocity, $L$ is the size of the largest eddies, and $\rho$ is the density of the fluid. Now, $\lambda$ is proportional to the inverse of the Kolmogorov time-scale $\tau_{\eta}$ [70], which is the turnover time of the smallest eddies in the flow. Using the definition $\tau_{\eta}=(\mu/\rho\epsilon)^{1/2}$, and the Taylor dissipation law $\epsilon\sim U^{3}/L$, one has $\tau_{\eta}=(\nu L)^{1/2}U^{-3/2}\sim(L/U)Re_{\lambda}^{-1}$ [72]. Usually, $L$ is set by the geometry of the system or the forcing mechanism, so that as one increases the forcing strength, $U$ will increase while $L$ will not. As a result, $\tau_{\eta}$ will decrease and $\lambda$ and $\mathrm{Wi}$ will increase; in addition, $Re_{\lambda}$ will increase. Our simulations, however, are performed at a single fixed value of $Re_{\lambda}$. So, strictly speaking, our simulations correspond to an experimental scenario in which $\mathrm{Wi}$ is increased, while keeping $Re_{\lambda}$ constant, by adjusting the forcing so that $U$ increases while $L$ decreases (leaving $UL$ unchanged). Nonetheless, we expect our results to hold to a good approximation even if $Re_{\lambda}$ increases along with $\mathrm{Wi}$. As $Re_{\lambda}$ increases, the distribution of strain-rates will become more heavy-tailed, i.e., the polymer will experience more extreme straining events [73, 74]. However, non-Gaussian extreme strain-rates have only a weak effect on polymer stretching; this result was shown for Rouse chains by [42] and is verified for chains with HI in Sec. V. So, our results on the effect of increasing $\mathrm{Wi}$ will be relevant even to situations in which the increase of $\mathrm{Wi}$ is accompanied by an increase in $Re_{\lambda}$.
As noted in Sec. II.1, our Brownian dynamics simulations assume a linear velocity field in the vicinity of the polymer which in turn requires $R_{\mathrm{m}}<\eta$ (where $\eta=(\nu^{3}/\epsilon)^{1/4}$ is the Kolmogorov length scale). This condition sets limits on the range of physical systems to which our results can be applied. For e.g., for a particular polymer and solvent combination, with given values of $R_{\mathrm{m}}$ and $\nu$, the requirement that $R_{\mathrm{m}}<\eta$ sets an upper limit on the range of dissipation rates $\epsilon$; of course, by varying the solvent viscosity one can alter the permissible range of $\epsilon$. Alternatively, if the solvent and dissipation rates are fixed, then the condition $R_{\mathrm{m}}<\eta$ limits the size of polymers to which our results may be applied (effectively limiting $R_{\mathrm{eq}}$ since our simulations use $R_{\mathrm{m}}^{2}/R_{\mathrm{eq}}^{2}=12000$).
@@ III Dumbbells and chains respond differently to HI
Figure 1: Mean rescaled extension $\langle R\rangle/R_{eq}$ for various values of the Weissenberg number $\mathrm{Wi}$, considering $h^{\star}$ = 0, 0.2, 0.49 [see the legend in panel (a)]. Results are presented for (a) dumbbells, (b) twenty-bead chains. (c) four-bead chains, and (d) ten-bead chains. The inset in panel (a) is a zoom of the main panel. The three curves in panel (c) almost entirely overlap.
On comparing the results without HI ($h^{\star}=0$) to those with HI ($h^{\star}=0.2$ and $0.49$), we see that HI marginally increases the stretching of dumbbells [Fig. 1(a)], while it strongly decreases the stretching of twenty-bead chains [Fig. 1(b)]. The effect of HI changes gradually from weak promotion to strong suppression of extension, as $N_{\mathrm{b}}$ is increased from 2 to 20, as evidenced by the results for $N_{\mathrm{b}}$ = 4 and 10 in Fig. 1(c) and Fig. 1(d), respectively. The slight enhancement of extension in case of dumbbells is in agreement with recent analytical results, derived using the Batchelor-Kraichnan model of turbulent transport [17].
@@
So why do HI affect dumbbells and multi-bead chains differently? A plausible explanation is provided by the mechanism of hydrodynamic shielding. At small $\mathrm{Wi}$, chains retract into coiled configurations in which the outer beads hydrodynamically shield the inner beads and thereby weaken the effect of fluid drag on the chain. As a consequence, small $\mathrm{Wi}$ chains stretch less in the presence of HI. Now, a dumbbell cannot form a coil and is by definition always straight. In this constrained arrangement, the effect of the disturbance flow created by the two beads is to increase the apparent relaxation time scale (as is analytically demonstrable [17] via the consistent-preaveraging approximation [75]). To check whether the influence of HI on multibead chains is associated with the formation of physical coils, we measure the hinge angles $\theta_{i}$ ($i=1,2,\dots,N_{\mathrm{b}}-2$) between successive links, and then compute the average angle $\overline{\theta}=\left(\sum_{1}^{N_{\mathrm{b}}-2}\theta_{i}\right)/(N_{\mathrm{b}}-2)$ for each chain. Note that a small end-to-end extension for a multibead chain could result from the links shrinking while the chain remains straight, or from the links retaining their length while the angles vary so that the chain folds into a physical coil. The difference between these two configurations will appear in the value of $\overline{\theta}$, which will be, respectively, near to and far from $0^{\circ}$ for straight and coiled chains. The PDF of $\overline{\theta}$ is presented for twenty-bead chains, for $\mathrm{Wi}=0.3$ and $0.5$, in Figs. 2(a-b). Increasing $h^{\star}$ increases the proportion of configurations with average inter-link angles close to $90^{\circ}$, which is consistent with HI acting to maintain chains in coiled configurations.